Use the ideas of additive and multiplicative inverses of rational numbers to solve the following equation:
1/3x-7=25

Found 2 solutions by rfer, mananth:
Answer by rfer(16322)

1/3x-7=25
1/3x=32
x=32/(1/3)
x=32*3/1
x=96

Answer by mananth(16946)

1/3x-7=25
..
add 7
1/3 *x -7+7=25+7
1/3*x=32
multiply by 3
3*x/3 =32*3
x=96
..